OpenStudy (anonymous):

If h(x)=(fog)(x) and h(x)=4(x+1)^2 find f(x) and g(x).

OpenStudy (amoodarya):

it has not a unique solution ,if we don't know about f(x) , or g(x) for example \[f(g(x))=4(x+1)^2\\f(x)=4x^2 ,g(x)=x+1\\ \space another \\f(x)=x^2 , g(x)=2x+2 \rightarrow f(2x+2)=(2x+2)^2=4 (x+1)^2\\-\space another -one\ \\f(x)=4x ,g(x)=(x+1)^2\\]

OpenStudy (amoodarya):

\[f(g(x))=4(x+1)^2\\f(x)=4x^2 ,g(x)=x+1\\ \space another \\f(x)=x^2 , g(x)=2x+2 \rightarrow f(2x+2)=(2x+2)^2=4 (x+1)^2\\-\space another -one\ \\f(x)=4x ,g(x)=(x+1)^2\\\]

OpenStudy (anonymous):

I'm confused. is the answer f(x)=4x2,g(x)=x+1 or f(x)=4x,g(x)=(x+1)2 ? I attached the answer choices. thanks!!

OpenStudy (amoodarya):

ok , it suffices to check all the choices

OpenStudy (amoodarya):

find it ?

OpenStudy (anonymous):

nope, i'm really confused still :/

OpenStudy (anonymous):

is it A?

OpenStudy (anonymous):

fml it's actually d if anyone is wondering.